(1 point) Book Problem 9 Find the volume of the solid obtained by rotating the region bounded by the curves: 12 6 x ; about y 3x , y = = Volume
(1 point) Book Problem 11 Find the volume of the solid obtained by rotating the region bounded by the curves: a2/4 22 ; about x =-3. y = x Volume:
Answer #1

Salution Roovi 36 4,12) ( 3 eaet kthod 2 d Racius 36 AC1= (12-y) 2.71 36 -( g2 49 ACy -f 3 3 36 Volumi ACUO 2 4 27 3 3 2. 3 6Salution feasgus Radiy 2. フ= V2 O1-3 ghel method Pocius 342 Kaignt= (3- 2+)( ACml 271 2 Valum ACm))dn 2 2 3m) -3n2 - 3 n 3 4